# The impact of noise on the simulation of NMR spectroscopy on NISQ devices

### v1

@@ Abstract
We present the simulation of nuclear magnetic resonance (NMR) spectroscopy of small organic molecules with two promising quantum computing platforms, namely IBM’s quantum processors based on superconducting qubits and IonQ’s Aria trapped ion quantum computer addressed via Amazon Bracket. We analyze the impact of noise on the obtained NMR spectra, and we formulate an effective decoherence rate that quantifies the threshold noise that our proposed algorithm can tolerate. Furthermore we showcase how our noise analysis allows us to improve the spectra. Our investigations pave the way to better employ such application-driven quantum tasks on current noisy quantum devices.
@@ I Introduction
The fragility of quantum coherence, owing to various sources of noise, is the main challenge ahead in unlocking the full potential of quantum technologies such as the realization of a fault tolerant quantum computer. To tackle such a challenge, various platforms have been proposed such as superconducting qubits [1, 2, 3], silicon spin qubits [4], photonics [5], neutral atoms [6], and trapped ions [7, 8]. Despite all the great advances, near-term quantum computers are prone to decoherence and other noisy events like crosstalk [9] and coherent overroation [10]. Scaling these systems to a regime beyond the capabilities of classical computers remains an open challenge [11]. To this end, various mitigation techniques [12, 13] have been introduced to increase the applicability of current Noisy Intermediate-Scale Quantum (NISQ) devices. However, the computational overhead of the implementation of these mitigation techniques is a crucial drawback.
@@
While it has been established that a fault-tolerant quantum computer has the potential to speed-up various computational problems [14] such as factoring large numbers [15, 16], finding applications for which NISQ devices can outperform classical computers has proven to be quite difficult. One of the most promising candidates for such applications is the digital quantum simulation of quantum matter [17] to address open questions in the field of quantum chemistry [18, 19] and condensed matter physics [20, 21]. In this respect, the variational quantum algorithms have been introduced [22], as hybrid quantum-classical approaches, which are quite attractive due to their tolerance to noise [23], but suffer from practical obstacles such as the “barren plateau” [24].
@@
In recent years there has been a growing interest in non-variational approaches to digital quantum simulation of quantum matter, including the Trotterized time evolution [25, 26, 27]. An important topic that fits very well to these non-variational approaches is the digital quantum simulation of nuclear magnetic resonance (NMR) spectroscopy [28], which has been studied using trapped-ion quantum computers [29]. The NMR problem addresses the time evolution of a many-body system that is an inherently computationally hard problem, due to the exponential growth of the computational space with the number of particles. While powerful classical methods exist that can approximate the NMR spectrum, they can reach their limits for approx. 50 spins and low magnetic fields. Therefore, advances in the digital quantum simulation with quantum processors would be a valuable breakthrough. However, while the Trotterized time evolution has a proven quantum advantage in the absence of noise [30], the impact of noise on the practicality of this approach is an active area of research [31]. As NMR spectroscopy plays a pivotal role in a wide range of industries from material chemistry [32, 33] to medical related bio-medicine [34, 35], it is a perfect example to explore the capabilities of NISQ devices.
@@
Figure 1: NMR spectrum of the non-exchangeable protons of cis-3-chloroacrylic acid simulated on NISQ devices. The solid orange curve depicts the simulation using IBM Perth, the dash-dotted blue curve is the simulation using IonQ Aria via Amazon Bracket, and the solid black curve shows the results from exact diagonalization. We have performed a time evolution for the spin Hamiltonian Eq. (2) at $B=11.7~{}T$ with Trotter step-size of $\Delta t=0.01$ and 81 and 61 Trotter-steps for IBM Perth and IonQ Aria via Amazon Bracket, respectively. For more details, see Sec. III.
@@
In this work, we fill the gap and provide insight into the noise tolerance of the digital quantum simulations of NMR experiments with NISQ devices. To this end, we present the simulation of NMR spectroscopy of small organic molecules on publicly available quantum processors, such as IBM’s quantum processors based on superconducting qubits and IonQ’s Aria trapped ion quantum computer available via Amazon Bracket, as it is showcased in Fig. 1 for NMR spectrum of cis-3-chloroacrylic acid. The simulation on quantum computers nicely reproduce the known spectrum. However, even for such a small molecule [see Eq. (2)], we observe the signatures of noise as slight deviations from the predicated resonances. This motivates further investigation of the impact of noise on quantum simulation of NMR spectroscopy, which is the main objective of the present work. In the actual NMR experiments, the nuclear spins unavoidably interact with their environment, which is similar to the noise acting on qubits. Therefore simulation of NMR spectra on quantum computers can endure a certain amount of noise. We quantify this amount of noise and try to find a threshold for the noise that can be tolerated.
@@ II The Spin Hamiltonian
The chemical shifts and the spin-spin couplings were determined to fit the experimental spectrum [40]. In addition, for 1,2,4-trichlorobenzene, the spin-spin couplings for the two opposing protons were taken from [41].
@@ III Method
In a 1D NMR experiment, an oscillatory external probe ($B^{\prime}(t)=B^{\prime}_{0}e^{i\omega t}$) is applied perpendicular to the static magnetic field mentioned above. From now on, we assume that the static magnetic field ($B=11.7~{}T$) is applied in the $x$ direction and the oscillatory one in the $z$ direction. Such a probe polarizes the nuclear spins and induces a time-dependent magnetic moment, that within linear response theory (linear order response in $B_{0}^{\prime}$), has the form
[…]
where $A(s)$ is the correlation function that can be formulated as
$A(t)=\frac{-i}{\hbar}\text{Tr}\{S^{z}_{\rm tot}(t)S^{z}_{\rm tot}\rho_{\rm eq}\},$ (6)
where $S^{z}_{\rm tot}=\sum_{i}S^{z}_{i}$ is the total nuclear spin that evolves in time as
$S^{z}_{\rm tot}(t)=e^{iHt}S^{z}_{\rm tot}e^{-iHt},$ (7)
and $\rho_{\rm eq}$ is the initial equilibrium density matrix (in the absence of the probe). As the temperature is quite large in the NMR experiment, compared to all energy scales in Eq. (1), we assume that all initial states are equally probable, resulting in a density operator of the form $\rho_{\rm eq}=\frac{1}{2^{N}}1\kern-2.5pt\text{l}$.
The NMR spectrum is measured based on the molecule absorption, which in turn reveals the internal structure of the molecule. The NMR spectrum is defined as
[…]
@@
Figure 2: Circuit corresponding to one Trotter step ($\Delta t=0.01$) of the Trotterized time evolution of the Hamiltonian Eq. (2) for IBM devices with native gates CNOT, $X$, $\sqrt{X}$, and virtual $R_{z}$. For the 2-spin molecule, the interaction can be decomposed into three CNOT gates in the second-order Trotter expansion.
In order to find an algorithm for obtaining such a correlation function on a quantum processor, we consider a Trotterized time evolution [42, 43], where the the Hamiltonian of the NMR problem is written in the sum of partial Hamiltonians $\hat{H}=\sum_{k}\hat{H}_{k}$ and the time evolution is implemented according to the Trotterization formula
$e^{-\textrm{i}\hat{H}t}\approx\left[\prod_{k}e^{-\textrm{i}\hat{H}_{k}\Delta t}\right]^{N}\,,$ (9)
with $\Delta t$ being the Trotter-step size and the approximation is controlled by the size of $\Delta t$ and $N$ [43]. For the NMR problem, the $\hat{H}_{k}$ terms correspond to the onsite energy terms and the interaction terms between spins. The time propagation under a partial Hamiltonian can be implemented on the quantum computer directly, using the unitary gates $\hat{U}$ on the quantum computer.
@@
In a Trotterized time evolution, the Trotter time step should be, on the one hand, small enough to minimize the Trotter error and adequately cover the entire spectrum of the molecule. On the other hand, the step size must be large enough to reduce the total number of Trotter steps required. In this respect, it is advantageous to perform the unitary transformation $H\to H-B\sum_{i}S^{x}_{i}$ that transforms the total magnetization in $z$-direction as
@@
Consequently, the Fourier transform of $\langle S^{z}_{\rm tot}(t)S^{z}_{\rm tot}(0)\rangle$ becomes
[…]
@@
the response functions can be rewritten according to
$\langle S^{z/y}_{\rm tot}(t)S^{z}_{\rm tot}(0)\rangle=2\sum_{m^{0}_{n}>0}m^{0}_{n}\langle m_{n}(t)|S^{z/y}_{\rm tot}|m_{n}(t)\rangle~{},$ (14)
@@
Figure 3: (a) Spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) from experimental NMR spectroscopy [40]. The marked peak at approx. 7.26 ppm is known to be caused by the solvent. (b) Comparison of the simulation with exact diagonalization, shown by the solid black curve, with noise-free simulation illustrated by the dashed blue curve. For the noise-free simulation, we have used the algorithm presented in Sec. III with 1000 Trotter steps of the size $\Delta t=0.005$.
In order to verify the outlined algorithm, we use a noise-free simulator to obtain the NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) molecule and we compare it to the known experimental results [40]. We choose the decoherence rate $\Gamma$, see Eq. (8), such that the simulated broadening of the peaks matches approximately the experimental broadening. As is shown in Fig. 3, the simulated results agree very well with the experimental data, apart from the marked peak that is known to be caused by the solvent in which the sample is embedded.
@@ IV Results on NISQ devices
In the following section, we present the outcome of executing the aforementioned method for organic molecules on quantum computing platforms. In the 1D NMR experiment, the NMR response is measured at sufficiently long times after the applied oscillatory magnetic field, such that the spin dynamics reach a steady state, in which the average magnetization is zero. However, in the performed simulation on current quantum computing platforms, we were restricted with regard to the total number of the circuits, i.e., the number of Trotter steps for the time evolution. We have observed that such a maximum number of Trotter steps is smaller than the spin relaxation times, and hence the oscillations in the averaged magnetization did not necessarily vanish at the end of the simulation, see Appendix A. Such a restriction also poses a maximum frequency grid for the spectrum. In order to be able to obtain the spectrum more finely, we have added additional zeros in the averaged magnetization, as is discussed in Appendix A.
@@
Figure 4: Comparison of the NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) simulated on IBM Perth on two different days, depicted as the orange and green curves, with the exact results shown as the black curve. For the simulations on IBM, we have performed a time evolution with 21 Trotter steps of the size $\Delta t=0.005$.
Fig. 4 depicts the spectrum of 1,2,4-trichlorobenzene simulated using the IBM Perth devise on two different days. As illustrated, the impact of noise on the simulations is so significant that we cannot reproduce all the fine structures of the eaxct NMR spectrum. This is due to the fact that the Hamiltonian of 1,2,4-trichlorobenzene has three spins with all-to-all interactions, see Eq. (4), On the IBM-Q device, we selected a three-qubit chain to execute the algorithm. Due to the Hamiltonian’s all-to-all interaction, it is necessary to swap the positions of the qubits. For simulating NMR spectra of 1,2,4-trichlorobenzene, we employed a second-order Trotter expansion. This approach results in application of 15 CNOT gates during each Trotter step. Moreover, due to this large-depth circuit, we have picked a smaller number of Trotter steps. This number is far below the required time for the average magnetization to reach zero, and hence the added zeros (see the above paragraph) cause an abrupt change in the magnetization. This, in turn, results in the oscillations around the center of the rotating frame.
@@ V Noise analysis
In this section, we analyze the impact of noise on the obtained NMR spectrum and construct a noise model which can reproduce most of the features of the shown results from the IBM quantum processors. We also define an effective decoherence rate that quantifies the amount of noise in a circuit, and the limitations it implies to the accuracy of the resulting spectrum.
V.1 Error sources
Disentangling the qubits from their embedded environments is one of the most challenging task of the realizations of quantum processors. At the same time modeling the noise that occurs in the quantum chips is quite complex. Typically each single qubit is characterized by the parameters $T_{1}$ and $T_{2}^{*}$ [46, 47]. The former is the damping time scale, i.e., the time scale after which a qubit in an excited state will decay to its ground state. The second time scale $T_{2}^{*}$ is pure dephasing time, which refers to the time scale, after which the qubit’s phase is lost (when removing the contribution from damping). Another source of error, apart from the imperfections of the qubits, is the erroneous unitary operation implemented by the gate. Such coherent error can be interpreted as effective additional unitary gate. Additionally, leakage out of the computational basis of the qubit [48], and a fast return back to mentioned basis may be mapped to incoherent errors.
V.2 Discrete noise model in quantum circuits
In order to account for the incoherent errors, Lindbladian master equations are used routinely. In this approach, the density operator of the qubits is described by the Lindblad master equation
Figure 5: (a) Comparison of the NMR spectrum of cis-3-chloroacrylic acid simulated on IBM Perth (solid orange curve) with on-gate noise model emulation (dashed blue curve) for 81 Trotter steps of size $\Delta t=0.01$. (b) Comparison of the NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) simulated on IBM Perth, shown as orange solid curve, with both the numerical model emulation including only depolarising noise shown as the dashed blue curve, and a noise model including also coherent errors illustrated by the dotted magenta curve, for 21 Trotter steps of size $\Delta t=0.005$. In the latter noise model, we have assumed a coherent Z-error with angle $\phi_{z}=-0.027$, after each gate operation. After each two-qubit gate operation (here, CNOT operations), we have assumed a coherent error of the form $R_{x}$, with angle $\phi_{x}=0.05,-0.047$ on the control and target qubits, respectively, and an additional error $R_{z}(-0.05)$ solely on the control qubit.
$\dot{\rho}=\mathcal{L}\rho$ $=-i[H_{\textrm{gates}}(t),\rho]$
$+\sum_{\alpha}\gamma_{\alpha}\left(O_{\alpha}\rho O_{\alpha}^{\dagger}-\frac{1}{2}\{O_{\alpha}O_{\alpha}^{\dagger},\rho\}\right),$ (15)
where the Hamiltonian $H_{\textrm{gates}}(t)$ describes coherent evolution of the qubits (the effect of gates) and the terms on the second line describe the incoherent noise. Here, qubit $i$ damping corresponds to operator $O_{i,-}=\sigma_{i}^{-}$ with a rate of $\gamma_{i,-}=\gamma^{i}_{\rm damp}=1/T_{1}$, and qubit pure dephasing corresponds to $O_{i,z}=\sigma^{i}_{z}$ with a rate of $2\gamma_{i,z}=\gamma^{i}_{\rm deph}=1/T_{2}^{*}$. Additionally, qubit depolarising noise corresponds to the summation over the three Pauli operators with factors $\gamma^{i}_{\rm depol}/4$.
The Lindblad superoperator also describes incoherent gate errors on the level of quantum circuits [31, 49]. Here, each ideal gate $\hat{U}$ is replaced by a gate-noise superoperators ${\cal N}{\cal U}$, where
${\cal N}=e^{t_{\textrm{gate}}\mathcal{L}_{N}}\,,$ (16)
and $t_{\textrm{gate}}$ is the gate time. The form of the noise Lindbladian $\mathcal{L}_{N}$ may be modified because it represents the effect of incoherent noise acting during the implementation of the gate. In particular, while the gates corresponding to small-angle rotations will keep the nature of the noise intact, the large-angle rotations will drastically modify the form of the noise, usually leading to a form close to depolarizing noise [49].
In our modeling of IBM’s quantum computing devices, we model the gate errors as depolarizing noise as all the native gates are corresponding to large angle rotations (except the $R_{z}$ which is a virtual gate). This model is consistent with noise spectroscopy performed on the IBM devices [50]. To fix the used strength of the depolarising noise after each gate, we use the universal connection between incoherent noise and gate fidelity [51], which yields
$\epsilon=\frac{3d}{4(d+1)}t_{\rm gate}\sum_{i}^{n}\gamma^{i}_{\textrm{depol}},$ (17)
where $n$ is the number of qubits the gate is acting on, and $d=2^{n}$ is the dimension of the space. The values for $\epsilon$ are obtained from IBM calibration data. We also account for the inherent noise during the qubit idling, i.e., considering damping and dephasing of qubits that are not involved in a two-qubit gates.
Figure 5(a) illustrates the comparison of the simulation which includes only depolarising noise according to Eq. (17) with the IBM results for cis-3-chloroacrylic acid. As it is shown the agreement is rather good. The peak positions show slight differences and also vary between different runs (see Appendix D), which is a manifestation of the effect of quasistatic coherent errors. Also the peak widths can vary between different runs.
Figure 5(b) shows the comparison of the simulation which includes only depolarising noise (blue dashed lines) with the IBM results for the case of 1,2,4-trichlorobenzene. This system involves more interaction terms and hence more gates in the circuit for time evolution [in total 15 CNOTs per Trotter step, in comparison to 3 CNOTs in case (a)]. As we see, here the broadenings of the peaks are not well reproduced by the simulation with including only depolarising noise. This difference is due to coherent errors, as demonstrated in the second numerical simulation (dotted lines), where we include coherent errors by applying an additional $R_{z}$ and $R_{x}$ on each involved gate-type determined by a fitting routine. We use angles whose magnitudes are consistent with the noise spectroscopy of Ref. [50] and correspond to gate errors that are smaller than given in the calibration data. (The size of the incoherent errors were kept unchanged). As illustrated, with the proper choice of angles corresponding to coherent errors, we are able to reproduce the main features of the simulated spectrum on IBM. This demonstrates that also coherent errors can effectively broaden the spectrum in large-depth circuits.
V.3 Effective decoherence rate
@@
Having constructed an effective noise model, in this section we aim to define an effective decoherence rate that quantifies the amount of noise in a circuit, and the limitations it implies to the accuracy of the resulting spectrum. The noise during the quantum computation corresponds to the simulation of an effective open quantum system [52]. Motivated by this, we define the effective decoherence rate as
$\Gamma_{\rm eff}=\frac{\sum_{j}^{N_{\rm gates}}\epsilon_{j}}{N\Delta t},$ (18)
with the summation performed over all gates applied at one Trotter circuit (including identity gates), $\Delta t$ being the simulated (Trotter) time step, $\epsilon_{j}$ error of the $j$-th applied gate, and $N$ the number of spins. It is important to note that the effective decoherence rate is inversely proportional to the Trotter step size [31]. This rate is expected to describe the error in the simulated energy-levels of the system and therefore also the broadening of peaks in the simulated spectra. To verify this hypothesis, we look at the simulated NMR spectrum on the IBMQ cloud for various Trotter steps. As is shown in Fig. 6, the broadening of the peaks in the simulated spectrum is in good agreement with the estimated effective broadenings, illustrated for various trotter steps. This indicates that indeed the impact of noise is less pronounced at larger trotter steps. However, there is an upper bound on the choice of the maximum Trotter step in order to keep the Trotterization errors small. The Trotter error is detectable in the simulation for $\Delta t=0.02$ (see Appendix C).

### v3

@@ Abstract
With the surge of quantum computing platforms that continue to push the boundaries of capabilities of noisy intermediate-scale quantum computers, there is a growing interest in finding relevant applications and quantifying the corresponding error budgets. We present a simulation of nuclear magnetic resonance (NMR) spectroscopy of small organic molecules on publicly available cloud quantum computers. We are using two quantum computing platforms, namely IBM’s quantum processors based on superconducting qubits and IonQ’s Aria trapped ion quantum computer addressed via Amazon Braket. We analyze the impact of noise on the obtained NMR spectra, and we formulate an effective decoherence rate that quantifies the threshold noise that our proposed algorithm can tolerate. We show that the effective decoherence rate can be calculated using simple fidelity metrics that are available by cloud quantum computing providers. Our investigation paves the way to better employ such application-driven quantum tasks on current noisy quantum devices.
@@ I Introduction
Among potential applications for quantum computers, it is clear that quantum simulation, the goal of solving intrinsically quantum mechanical problems, is one of the most interesting and promising ones. Digital quantum simulation  [1] is often considered to address open questions in the field of quantum chemistry [2, 3] and condensed matter physics [4, 5]. In this respect, the variational quantum algorithms have been introduced [6], as a mean to find the low-temperature properties, i.e., ground and low lying states, but suffer from practical obstacles such as the “barren plateau”  [7].
@@
In recent years there has been a growing interest in non-variational approaches to digital quantum simulation of quantum matter, including the Trotterized time evolution [8, 9, 10]. The advantage of such approaches over variational algorithms is the proven exponential quantum advantage when simulating a time evolution. An important topic that requires a quantum mechanical time evolutions is simulation of the nuclear magnetic resonance (NMR) spectroscopy [11, 12], that has been recently studied using trapped-ion quantum computers [13]. The NMR problem addresses the time evolution of a many-body system that is inherently computationally hard [14, 15], due to the exponential growth of the computational space with the number of particles. While powerful classical methods exist that can approximate the NMR spectrum, they can reach their limits for approx. 50 spins and low magnetic fields. Therefore, advances in the digital quantum simulation with quantum processors would be a valuable breakthrough.
Given that simulating NMR spectrum seems to be an ideal problem for quantum computers, it is of course also an interesting question to ask if it is feasible to be tackled using the current generation of noisy intermediate scale quantum (NISQ) computers. While it has been established that a fault-tolerant quantum computer has the potential to speed-up various computational problems [16] such as factoring large numbers [17, 18], finding applications for which NISQ devices can outperform classical computers has proven to be quite difficult. The fragility of quantum coherence, owing to various sources of noise, is the main challenge ahead in unlocking the full potential of NISQ computers. To tackle such a challenge, various platforms have been proposed such as superconducting qubits [19, 20, 21], silicon spin qubits [22], photonics [23], neutral atoms [24], and trapped ions [25, 26]. Despite all the great advances, near-term quantum computers are prone to decoherence and other noisy events like crosstalk [27] and coherent overroation [28]. Scaling these systems to a regime beyond the capabilities of classical computers remains an open challenge [29]. To this end, various mitigation techniques [30, 31] have been introduced to increase the applicability of current Noisy Intermediate-Scale Quantum (NISQ) devices. However, the computational overhead of the implementation of these mitigation techniques is a crucial drawback.
The impact of noise on the practicality of simulating a Trotterized time evolution is an active area of research [32]. As NMR spectroscopy plays a pivotal role in a wide range of industries from material chemistry [33, 34] to medical related bio-medicine [35, 36], it is a perfect example to explore the capabilities of NISQ devices. In this work, we fill the gap and provide insight into the noise tolerance of the digital quantum simulations of NMR experiments with NISQ devices. A key goal here is also the use of publicly availble devices that are quite easily available to the broader community. This will open the path to further improvements that might achieve practical applications in NMR spectroscopy.
@@
Figure 1: NMR spectrum of the non-exchangeable protons of cis-3-chloroacrylic acid simulated on NISQ devices. The solid orange curve depicts the simulation using IBM Perth, the dash-dotted blue curve is the simulation using IonQ Aria via Amazon Braket, and the solid black curve shows the results from exact diagonalization. We have performed a time evolution for the spin Hamiltonian Eq. (2) at B=11.7TB=11.7~T with Trotter step-size of $\tau=0.01$ and 81 and 61 Trotter-steps for IBM Perth and IonQ Aria via Amazon Braket, respectively.
@@
To this end, we present the simulation of NMR spectroscopy of small organic molecules on publicly available quantum processors, such as IBM’s quantum processors based on superconducting qubits and IonQ’s Aria trapped ion quantum computer available via Amazon Braket, as it is showcased in Fig. 1 for NMR spectrum of cis-3-chloroacrylic acid. The simulation on quantum computers nicely reproduce the known spectrum. However, even for such a small molecule [see Eq. (2)], we observe the signatures of noise as broadening of the predicted resonances. This motivates further investigation of the impact of noise on quantum simulation of NMR spectroscopy, which is the main objective of the present work. In the actual NMR experiments, the nuclear spins unavoidably interact with their environment, which is similar to the noise acting on qubits. Therefore simulation of NMR spectra on quantum computers can endure a certain amount of noise. We quantify this amount of noise and try to find a threshold for the noise that can be tolerated.
@@ II The Spin Hamiltonian
The chemical shifts and the spin-spin couplings were determined to fit the experimental spectrum [41]. In addition, for 1,2,4-trichlorobenzene, the spin-spin couplings for the two opposing protons were taken from Ref. [42].
@@ III Method
In a 1D NMR experiment, an oscillatory external probe is applied perpendicular to the static magnetic field mentioned above. From now on, we assume that the static magnetic field (B=11.7TB=11.7~T) is applied in the $x$ direction. Such a probe polarizes the nuclear spins and induces a time-dependent magnetic moment that can be expressed in terms of correlation function [43] $A(t)=\text{Tr}\{S^{z}_{\rm tot}(t)S^{z}_{\rm tot}\}$ where $S^{z}_{\rm tot}=\sum_{i}S^{z}_{i}$ is the total nuclear spin that evolves in time as
$S^{z}_{\rm tot}(t)=e^{iHt}S^{z}_{\rm tot}e^{-iHt}.$ (5)
The NMR spectrum can be expressed in terms of the Fourier transform of the mentioned correlation function
$A(\omega)=\text{Re}\left\{\int_{0}^{\infty}~dt~e^{-i\omega t-\Gamma t}\text{Tr}\{S^{z}_{\rm tot}(t)S^{z}_{\rm tot}\}\right\},$ (6)
@@
Figure 2: Circuit corresponding to one Trotter step ($\tau=0.01$) of the Trotterized time evolution of the Hamiltonian Eq. (2) for IBM devices with native gates CNOT, $X$, $\sqrt{X}$, and virtual $R_{z}$. For the 2-spin molecule, the interaction can be decomposed into three CNOT gates in the second-order Trotter expansion.
In order to find an algorithm for obtaining such a correlation function on a quantum processor, we consider a Trotterized time evolution [44, 45], where the Hamiltonian of the NMR problem is written as the sum of partial Hamiltonians $\hat{H}=\sum_{k}\hat{H}_{k}$ and the time evolution is implemented according to the Trotterization formula
$e^{-\textrm{i}\hat{H}t}\approx\left[\prod_{k}e^{-\textrm{i}\hat{H}_{k}\tau}\right]^{N}\,$, (7)
with $\tau$ being the Trotter-step size and the approximation is controlled by the size of $\tau$ and $N$ [45]. For the NMR problem, the $\hat{H}_{k}$ terms correspond to the onsite energy terms and the interaction terms between spins. The time propagation under a partial Hamiltonian can be implemented on the quantum computer directly, using the unitary gates $\hat{U}$ on the quantum computer.
@@
In a Trotterized time evolution, the Trotter time step should be, on the one hand, small enough to minimize the Trotter error and adequately cover the entire spectrum of the molecule. On the other hand, the step size must be large enough to reduce the total number of Trotter steps required to simulate a sufficient length of the time evolution. As we will see later, the size of the trotter step is inversely related to the effective width of the peaks in the simulated NMR spectrum. In this respect, it is advantageous to perform the unitary transformation $H\to H-B\sum_{i}S^{x}_{i}$ that transforms the total magnetization in $z$-direction as
@@
Consequently, we get (after neglecting counter-rotating terms)
$A(\omega)=\text{Re}\left\{\int~dt~e^{-i(\omega-B)t-\Gamma t}\text{Tr}\{S^{z}_{\rm tot}(t)S^{z}_{\rm tot}(0)\}\right\}$
$-\text{Im}\left\{\int~dt~e^{-i(\omega-B)t-\Gamma t}\text{Tr}\{S^{y}_{\rm tot}(t)S^{z}_{\rm tot}(0)\}\right\}.$ (10)
@@
the correlation functions can be rewritten according to
$\text{Tr}\{S^{z/y}_{\rm tot}(t)S^{z}_{\rm tot}(0)\}=2\sum_{m^{0}_{n}>0}m^{0}_{n}\langle m_{n}(t)|S^{z/y}_{\rm tot}|m_{n}(t)\rangle~,$ (12)
@@
Figure 3: (a) Spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) from experimental NMR spectroscopy [41]. The marked peak at approx. 7.26 ppm is known to be caused by the solvent. (b) Comparison of the simulation with exact diagonalization, shown by the solid black curve, with noise-free simulation illustrated by the dashed blue curve. For the noise-free simulation, we have used the algorithm presented in the Method section with 1000 Trotter steps of the size $\tau=0.005$.
In order to verify the outlined algorithm, we use a noise-free simulator to obtain the NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) molecule and we compare it to the known experimental results [41]. We choose the decoherence rate $\Gamma$, such that the simulated broadening of the peaks matches approximately the experimental broadening. As is shown in Fig. 3, the simulated results agree very well with the experimental data, apart from the marked peak that is known to be caused by the solvent in which the sample is embedded.
@@ IV Results on NISQ devices
In the following section, we present the outcome of executing the aforementioned method for organic molecules on quantum computing platforms. In the 1D NMR experiment, the NMR response is measured at sufficiently long times after the applied oscillatory magnetic field, such that the spin dynamics reach a steady state, in which the average magnetization is zero. However, in our runs on current quantum computing platforms, we were restricted with regard to the total depth of the circuits, i.e., the number of Trotter steps for the time evolution. We have observed that such a maximum number of Trotter steps is smaller than the spin relaxation times, and hence the oscillations in the averaged magnetization did not necessarily vanish at the end of the simulation, see Appendix A. Such a restriction also poses a maximum frequency grid for the spectrum. In order to be able to obtain the spectrum more finely, we have added additional zeros in the averaged magnetization.
@@
Figure 4: Comparison of the NMR spectrum of 1,2,4-trichlorobenzene ($C_{6}H_{3}Cl_{3}$) simulated on IBM Perth on two different days, depicted as the orange and green curves, with the exact results shown as the black curve. For the simulations on IBM, we have performed a time evolution with 21 Trotter steps of the size $\tau=0.005$.
Figure 4 depicts the spectrum of 1,2,4-trichlorobenzene simulated using the IBM Perth devise on two different days. As illustrated, the impact of noise on the simulations is so significant that we cannot reproduce all the fine structures of the eaxct NMR spectrum. This is due to the fact that the Hamiltonian of 1,2,4-trichlorobenzene has three spins with all-to-all interactions, see Eq. (4), On the IBM-Q device, we selected a three-qubit chain to execute the algorithm. Due to the Hamiltonian’s all-to-all interaction, it is necessary to swap the positions of the qubits. For simulating NMR spectra of 1,2,4-trichlorobenzene, we employed a second-order Trotter expansion. This approach results in application of 15 CNOT gates during each Trotter step. Moreover, due to this large-depth circuit, we have picked a smaller number of Trotter steps. This number is far below the required time for the average magnetization to reach zero, and hence the added zeros (see the above paragraph) cause an abrupt change in the magnetization. This, in turn, results in the oscillations around the center of the rotating frame.
@@ V Noise analysis
A critical component of every execution of any algorithm on current quantum computers is the effect of noise. For any algorithm, it would be of great use to have a way to construct estimates for the quality of the execution on a NISQ computer purely from the noise characteristics of the quantum gates. Normally, the noise characteristics are characterized by the quantum computing providers in terms of error probabilities $\epsilon_{kl}$ or gate fidelities $F_{kl}=1-\epsilon_{kl}$, where every gate in the implementation of a Trotterized time evolution, see Eq. 8, is associated with a Fidelity and error probability. We will now discuss why the width of the peaks in the simulated NMR spectra can be roughly estimated to be
$\Gamma_{\rm eff}=\frac{\sum_{kl}\epsilon_{kl}}{N\tau},$ (13)
where $\epsilon_{kl}$ is the error probability of each gate that is applied in the algorithm, see Eq. 8 and Fig. 2, where $N$ is the number of qubits in the simulation and $\tau$ is the Trotter time step as discussed above. In Fig. 5, we illustrate the simulated NMR spectrum on the IBM quantum computer for various Trotter steps. The figure also shows the estimate of the peak width $\Gamma_{\rm eff}$ which was purely done using data provided by IBM, and without any additional fitting. The results indicates that the scaling on the trotter step size behaves exactly as predicted by Eq. 13.
@@
The inverse proportion of the peak width with the Trotter time step is in itself a major result of this work. In principle, it is clear that the noise will simply scale linearly with the number of applied gates and the number of applied gates scales linearly with the Trotter time step if total simulation time is kept constant. But it is often not appreciated that this trade off exists and a larger Trotter step might be preferable (due to suppression of the impact of noise) despite the increased Trotter error.